HanAkoSolo (jamierox4ev3r):

Find all values of x such that sin 2x = sin x and ________________.(List the answers in increasing order.)

HanAkoSolo (jamierox4ev3r):

@ganeshie8

HanAkoSolo (jamierox4ev3r):

wow, hold on, a certain part of the question isn't coming out right...

HanAkoSolo (jamierox4ev3r):

"Find all values of x such that sin2x=sinx and \(0\le x \le 2\pi\). (List the answers in increasing order)

ganeshie8 (ganeshie8):

there are at least two ways to solve this il show you the first/fast way quick

ganeshie8 (ganeshie8):

\(\sin(2x) = \sin (x)\) since \(\sin(t)\) has a period of \(2\pi\), we must have \[2x=x+2n\pi \implies x = 2n\pi\tag{1}\] since \(\sin(t)=\sin(\pi-t)\), we must have \[2x=\pi-x+2n\pi\implies x =\dfrac{\pi}{3}+\dfrac{2n\pi}{3} \tag{2}\] plugin \(n=0,1,2\ldots\) and take the solutions that are within the given interval

ganeshie8 (ganeshie8):

if you don't like that method, try using below identity : \[\sin(2x) = 2\sin x\cos x\]

HanAkoSolo (jamierox4ev3r):

oh right! that's one of the double angle identities, if I'm not mistaken

ganeshie8 (ganeshie8):

Yes put everything on one side and try factoring

ganeshie8 (ganeshie8):

\(\sin 2x = \sin x\) \(2\sin x\cos x = \sin x\) \(\sin x(2\cos x-1) = 0\) use ur favorite zero product property

HanAkoSolo (jamierox4ev3r):

put everything on one side? So do you mean set this equation to zero?

ganeshie8 (ganeshie8):

yes, i did t that already for you

HanAkoSolo (jamierox4ev3r):

yep, that's what you meant. and oooh yeah you mentioned this before. My favorite indeed/

HanAkoSolo (jamierox4ev3r):

hmmm.... so upon solving, I see that we can have \(\Large 0, \frac{\pi}{3}, \pi , \frac{5\pi}{3}\)

HanAkoSolo (jamierox4ev3r):

Am i missing any?

ganeshie8 (ganeshie8):

you're missing just one

HanAkoSolo (jamierox4ev3r):

oh, would that be \(2\pi\)?

ganeshie8 (ganeshie8):

Yes, \(x=2\pi\) satisfies \(\sin(x)=0\)

HanAkoSolo (jamierox4ev3r):

I forgot that in this scenario, it wasn't excluded. Well alright, I think I got it.

ganeshie8 (ganeshie8):

good job!

HanAkoSolo (jamierox4ev3r):

Thank you! You're terrific at explaining things \(\ddot\smile\)

ganeshie8 (ganeshie8):

np im as terrific as you're smart at picking up on these :)